pemutations & combications from this how we have to solve the 45C 39 ?

Grade:11

1 Answers

Rishabh Kumar Mahato
7 years ago
Use the formula:nCr= n!/r!(n-r)!Where ! Means continuous consecutive multiplication...I.e n!=n(n-1)(n-2)(n-3)........3.2.1So 45C39:= 45!/39!6!The answer would be too big so it is left as it is.Though answer is:=45!/39!(45−39)!)=8.14506E+6= 8145060